# A Survey on LLM-as-a-Judge

### v3

@@ Abstract
Accurate and consistent evaluation is crucial for decision-making across numerous fields, yet it remains a challenging task due to inherent subjectivity, variability, and scale. Large Language Models (LLMs) have achieved remarkable success across diverse domains, leading to the emergence of ”LLM-as-a-Judge,” where LLMs are employed as evaluators for complex tasks. With their ability to process diverse data types and provide scalable, cost-effective, and consistent assessments, LLMs present a compelling alternative to traditional expert-driven evaluations. However, ensuring the reliability of LLM-as-a-Judge systems remains a significant challenge that requires careful design and standardization. This paper provides a comprehensive survey of LLM-as-a-Judge, addressing the core question: How can reliable LLM-as-a-Judge systems be built? We explore strategies to enhance reliability, including improving consistency, mitigating biases, and adapting to diverse assessment scenarios. Additionally, we propose methodologies for evaluating the reliability of LLM-as-a-Judge systems, supported by a novel benchmark designed for this purpose. To advance the development and real-world deployment of LLM-as-a-Judge systems, we also discussed practical applications, challenges, and future directions. This survey serves as a foundational reference for researchers and practitioners in this rapidly evolving field. The associated resources can be accessed at https://github.com/IDEA-FinAI/LLM-as-a-Judge.
@@ 1. Introduction
Before the era of LLMs, achieving a balance between comprehensive and scalable evaluation had posed a persistent challenge. On the one hand, commonly used subjective methods like expert-driven assessments (Shi et al., 2024; Gao et al., 2023a) integrate holistic reasoning and nuanced contextual understanding, making them the gold standard in comprehensiveness. However, these approaches are expensive, challenging to scale, and susceptible to inconsistency. On the other hand, objective assessment methods like automatic metrics offer strong scalability and consistency. For example, tools such as BLEU (Papineni et al., 2002) or ROUGE (Lin, 2004) can quickly score machine-generated translations or summaries against reference texts without human intervention. Yet these metrics, which heavily rely on surface-level lexical overlaps, often fail to capture deeper nuances, leading to poor performance in tasks like story generation or instructional texts (Schluter, 2017). As a solution to this persistent dilemma, “LLM-as-a-Judge” has emerged as a promising idea to combine the strengths of the above two evaluation methods, merging the scalability of automatic methods with the nuanced, context-sensitive reasoning found in expert judgments (Zheng et al., 2023b; Wang et al., 2023d; Zhu et al., 2023a; Li et al., 2023b; Chen et al., 2024e). Moreover, LLMs might become flexible enough to handle multimodal inputs (Chen et al., 2024d) under decent prompt learning or fine-tuning (Khattak et al., 2023). These advantages suggest that the LLM-as-a-Judge approach could serve as a novel and broadly applicable paradigm for addressing complex and open-ended evaluation problems.
While LLM-as-a-Judge holds significant potential as a scalable and adaptable evaluation framework compared to expert-driven human evaluations and traditional automated metrics (Wang et al., 2024a), its adoption is hindered by two key challenges. The first challenge is the absence of a systematic review, which reflects the lack of formal definitions, fragmented understanding, and casual usage practices in the study of LLM-as-a-Judge. These gaps make it difficult for researchers and practitioners to fully understand LLM-as-a-Judge and apply it effectively. Building on this, the second challenge involves addressing ongoing concerns about reliability (Yu et al., 2024b), as merely employing LLM-as-a-Judge does not guarantee evaluations that are both accurate and aligned with established standards. These challenges demand not only a deeper assessment of the outputs generated by LLM-as-a-Judge but also a crucial investigation into the question: How to build reliable LLM-as-a-Judge systems?
@@
The rest of this survey is organized as Figure 1. Section 2 provides an overview of the LLM-as-a-Judge field, including its definitions and categorization of existing methods. For a quick guide on the implementation of an LLM as a judge for specific scenarios, you can find answers in Quick Practice (2.5). Strategies for enhancing and evaluating the reliability of LLM-as-a-Judge systems are discussed in Sections 3,  4, and  5, respectively. Section 6 explores practical applications, while Sections 7 and 8 address challenges and outline future research directions. Finally, Section 9 presents our conclusions.
@@ 2. Background and Method
The ability of LLMs to mimic human reasoning and evaluate specific inputs against a set of predefined rules has paved the way for ”LLM-as-a-Judge.” Their scalability, adaptability, and cost-effectiveness make them well-suited for managing a growing number of evaluative tasks that were traditionally done by humans. These abilities are key in utilizing LLMs flexibly across various evaluation scenarios and objectives. The adoption of LLM-as-a-Judge has progressed rapidly. Initially, the primary focus of LLMs was on language generation and comprehension. With advancements in training paradigms like Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022a), LLMs became more aligned with human values and reasoning processes. This alignment has allowed LLMs to transition from generative tasks to evaluative roles. At its core, LLM-as-a-Judge refers to the use of LLMs to evaluate objects, actions, or decisions based on predefined rules, criteria, or preferences. It encompasses a broad spectrum of roles, including: Graders (Trung et al., 2024; Dong et al., 2023), Evaluators/Assessors (Li et al., 2024b; Zhang et al., 2024b), Critics (Ke et al., 2024; Xiong et al., 2024; Putta et al., 2024), Verifiers (Ling et al., 2024; Shinn et al., 2023; Wang et al., 2024c), Examiners (Bai et al., 2023a), Reward/Ranking Models (Yang et al., 2024b; Sun et al., 2023a; Luo et al., 2023a; Yuan et al., 2023), etc.
@@ 2.1.1. Generating scores
Figure 3. The template for Likert scale scoring from Gao et al. (2023a).
@@ 2.1.3. Conducting pairwise comparisons
Pairwise comparison is a well-established method that has significantly impacted a variety of fields (Qin et al., 2024). As noted by (Liu et al., 2024b), LLM and human evaluations are more aligned in the context of pairwise comparisons compared to score-based assessments. Numerous studies have demonstrated that pairwise comparative assessments outperform other judging methods in terms of positional consistency (Zheng et al., 2023c; Liusie et al., 2024). Furthermore, pairwise comparisons can be extended to more complex relation-based assessment frameworks, such as list-wise comparisons, using advanced ranking algorithms (Qin et al., 2024; Liu et al., 2024b), data filtering (Yuan et al., 2023). In pairwise comparative assessments, LLM-as-a-Judge is prompted to select the response that better answers the question at hand. To accommodate the possibility of a tie, several option modes are introduced. The Two-Option mode requires judges to choose the better response from two given options. The Three-Option mode introduces an additional choice, allowing judges to indicate a tie if neither response is preferable. Evaluations typically involve determining the outcomes of win, tie, or loss for responses (Wang et al., 2023d) through pairwise comparisons, with win rounds counted for each response. The Four-Option mode further expands the choices, allowing judges to classify responses as either a ”both good tie” or a ”both bad tie.”
@@
Figure 5. The template for pairwise comparison from Gao et al. (2023a)
@@ 2.2.2. Fine-tuned LLM
However, relying on external API for evaluation may introduce consideration about privacy leakage, and the opacity of API models also challenges the evaluation reproducibility. Therefore, subsequent studies recommend refining language models tailored for evaluations by emphasizing the use of pairwise comparisons or grading. For instance, PandaLM (Wang et al., 2023d) constructs data based on Alpaca instructions and GPT-3.5 annotation, and then fine-tunes LLaMA-7B (Touvron et al., 2023a) as an evaluator model. JudgeLM (Zhu et al., 2023a) constructs data from diversified instruction sets and GPT-4 annotations, and fine-tunes Vicuna (Touvron et al., 2023b) as a scalable evaluator model. Auto-J (Li et al., 2023b) constructs evaluation data upon multiple scenarios to train a generative evaluator model, which can provide both evaluation and critical opinion. Prometheus (Kim et al., 2023) defines thousands of evaluation criteria and construct a feedback dataset based on GPT-4, and fine-tunes a fine-grained evaluator model.
The typical process for fine-tuning a judge model involves three main steps. Step 1: Data Collection. The training data generally consists of three components: instructions, the objects to be evaluated, and evaluations. Instructions are typically sourced from instruction datasets, while evaluations can come from either GPT-4 or human annotations. Step 2-Prompt Design. The structure of the prompt template can vary based on the evaluation scheme, which already detailed in § 2.1. Step 3: Model Fine-Tuning. Using the designed prompts and collected data, the fine-tuning process for the evaluator model typically adheres to the instruction fine-tuning paradigm (Ouyang et al., 2022b). The model receives an instruction along with one or more responses to generate output that includes evaluation results and possibly explanations.
@@ 2.3.2. Normalizing the output logits
In addition, Self-evaluation (Hao et al., 2023) is also common using this method for LLM-as-a-Judge. It can be helpful to let the LLM evaluate itself by asking, ”Is this reasoning step correct?” and then reward it based on the probability of the next word being ”Yes.”
@@ 2.3.3. Selecting sentences
In addition to selecting specific tokens and normalizing the output logits, the content extracted by LLM-as-a-Judge may also be a sentence or paragraph. As showed in Figure 2, agent for reasoning task (Hao et al., 2023), builds a reasoning tree by iteratively considering the most promising reasoning steps (actions, sub-questions) by LLM-as-a-Judge.
@@ 2.4.1. LLM-as-a-Judge for Models
It is universally known that the best way to evaluate LLMs is human judgment, but collecting human annotations can be costly, time-consuming, and laborious (Ouyang et al., 2022b; Zheng et al., 2023d). Using strong LLMs (usually closed-source ones, e.g., GPT-4, Claude, ChatGPT) as an automated proxy for assessing LLMs has become a natural choice (Zhou et al., 2023). With appropriate prompt design, the quality of evaluation and agreement to human judgment can be promising (Dubois et al., 2023; Zheng et al., 2023d; Zhang et al., 2023c; Wang et al., 2023b). However, the cost concern still exists when calling the APIs of these proprietary models, especially when there is a frequent need for model validation on large-scale data. Moreover, closed-source LLM-as-a-Judge leads to low reproducibility due to potential changes in models behind the API. Some recent works have started to make attempts for open-source alternatives. SelFee (Ye et al., 2023) collects generations, feedback, and revised generations from ChatGPT and fine-tunes LLaMA models to build a critique model. Shepherd (Wang et al., 2023c) trains a model that can output critiques for single-response with the data of feedback from online communities and human annotation. PandaLM (Wang et al., 2023d) trains a model to conduct pairwise comparison for LLM Instruction Tuning Optimization, and Zheng et al. (2023d) also fine-tune Vicuna (Touvron et al., 2023b) on a 20K pairwise comparison dataset to explore the potential of open-source models as a more cost-friendly proxy.
@@ 2.4.2. LLM-as-a-Judge for Data
However, solely relying on LLM-as-a-Judge for data annotation poses challenges, particularly as the value of annotated data diminishes with the rapid improvement of model performance. To address this, approaches like Self-Taught Evaluator (Wang et al., 2024a) offer a promising alternative by eliminating the need for human annotations. This method leverages synthetic training data, starting with unlabeled instructions and generating contrasting outputs from models. These outputs are then used to train an LLM-as-a-Judge to produce reasoning traces and final judgments. With each iteration, the evaluator improves by learning from its refined predictions, creating a cycle of continuous self-enhancement. This iterative approach not only keeps annotations relevant but also ensures that evaluators evolve alongside advancing models.
@@ 2.4.3. LLM-as-a-Judge for Agents
There are two ways to apply LLM-as-a-Judge for an agent. One is to evaluate the entire process of the intelligent agent (Zhuge et al., 2024), and the other is to evaluate it at a specific stage in the agent framework process (Hao et al., 2023; Shinn et al., 2023). Both approaches are briefly illustrated in Figure 8. Using LLM as the brain of agent, an agentic system (Zhuge et al., 2024) could evaluate like a human, it would reduce the need for human involvement and eliminate the trade-off between thoroughness and effort. In addition, the agent (Shinn et al., 2023) can interact with the environment through language and receive feedback on actions through LLM to make decisions for the next action.

### v4

@@ Abstract
Accurate and consistent evaluation is crucial for decision-making across numerous fields, yet it remains a challenging task due to inherent subjectivity, variability, and scale. Large Language Models (LLMs) have achieved remarkable success across diverse domains, leading to the emergence of ”LLM-as-a-Judge,” where LLMs are employed as evaluators for complex tasks. With their ability to process diverse data types and provide scalable and flexible assessments, LLMs present a compelling alternative to traditional expert-driven evaluations. However, ensuring the reliability of LLM-as-a-Judge systems remains a significant challenge that requires careful design and standardization. This paper provides a comprehensive survey on LLM-as-a-Judge, offering a formal definition and a detailed classification of its use cases, while focusing on addressing the core question: How can reliable LLM-as-a-Judge systems be built? We explore strategies to enhance reliability, including improving consistency, mitigating biases, and adapting to diverse assessment scenarios. Additionally, we propose methodologies for evaluating the reliability of LLM-as-a-Judge systems, supported by a novel benchmark designed for this purpose. To advance the development and real-world deployment of LLM-as-a-Judge systems, we also discussed practical applications, challenges, and future directions. This survey serves as a foundational reference for researchers and practitioners in this rapidly evolving field. The associated resources can be accessed at https://awesome-llm-as-a-judge.github.io/.
@@ 1. Introduction
Before the era of LLMs, achieving a balance between comprehensive and scalable evaluation had posed a persistent challenge. On the one hand, commonly used subjective methods like expert-driven assessments (Shi et al., 2024; Gao et al., 2023b) integrate holistic reasoning and nuanced contextual understanding, making them the gold standard in comprehensiveness. However, these approaches are expensive, challenging to scale, and susceptible to inconsistency. On the other hand, objective assessment methods like automatic metrics offer strong scalability and consistency. For example, tools such as BLEU (Papineni et al., 2002) or ROUGE (Lin, 2004) can quickly score machine-generated translations or summaries against reference texts without human intervention. Yet these metrics, which heavily rely on surface-level lexical overlaps, often fail to capture deeper nuances, leading to poor performance in tasks like story generation or instructional texts (Schluter, 2017). As a solution to this persistent dilemma, “LLM-as-a-Judge” has emerged as a promising idea to combine the strengths of the above two evaluation methods, merging the scalability of automatic methods with the nuanced, context-sensitive reasoning found in expert judgments (Zheng et al., 2023b; Wang et al., 2023d; Zhu et al., 2023a; Li et al., 2023b; Chen et al., 2024e). Moreover, LLMs might become flexible enough to handle multimodal inputs (Chen et al., 2024d) under decent prompt learning or fine-tuning (Khattak et al., 2023). These advantages suggest that the LLM-as-a-Judge approach could serve as a novel and broadly applicable paradigm for addressing complex and open-ended evaluation problems.
While LLM-as-a-Judge holds significant potential as a scalable and adaptable evaluation framework compared to expert-driven human evaluations and traditional automated metrics (Wang et al., 2024b), its adoption is hindered by two key challenges. The first challenge is the absence of a systematic review, which reflects the lack of formal definitions, fragmented understanding, and casual usage practices in the study of LLM-as-a-Judge. These gaps make it difficult for researchers and practitioners to fully understand LLM-as-a-Judge and apply it effectively. Building on this, the second challenge involves addressing ongoing concerns about reliability (Yu et al., 2024b), as merely employing LLM-as-a-Judge does not guarantee evaluations that are both accurate and aligned with established standards. These challenges demand not only a deeper assessment of the outputs generated by LLM-as-a-Judge but also a crucial investigation into the question: How to build reliable LLM-as-a-Judge systems?
@@
The rest of this survey is organized as Figure 1. Section 2 provides an overview of the LLM-as-a-Judge field, including its definitions and categorization of existing methods. For a quick guide on the implementation of an LLM as a judge for specific scenarios, you can find answers in Quick Practice (2.5). Strategies for enhancing and evaluating the reliability of LLM-as-a-Judge systems are discussed in Sections 3,  4, and  5, respectively. Notably, in Section 6, we discuss the synergy between LLM-as-a-Judge and o1-like reasoning enhancement, where dynamic feedback is used to optimize reasoning paths and significantly improve the model’s ability to solve complex problems. Section 7 explores practical applications, while Sections 8 and 9 address challenges and outline future research directions. Finally, Section 10 presents our conclusions.
@@ 2. Background and Method
The ability of LLMs to mimic human reasoning and evaluate specific inputs against a set of predefined rules has paved the way for ”LLM-as-a-Judge.” Their scalability, adaptability, and cost-effectiveness make them well-suited for managing a growing number of evaluative tasks that were traditionally done by humans. These abilities are key in utilizing LLMs flexibly across various evaluation scenarios and objectives. The adoption of LLM-as-a-Judge has progressed rapidly. Initially, the primary focus of LLMs was on language generation and comprehension. With advancements in training paradigms like Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022a), LLMs became more aligned with human values and reasoning processes. This alignment has allowed LLMs to transition from generative tasks to evaluative roles. At its core, LLM-as-a-Judge refers to the use of LLMs to evaluate objects, actions, or decisions based on predefined rules, criteria, or preferences. It encompasses a broad spectrum of roles, including: Graders (Trung et al., 2024; Dong et al., 2023), Evaluators/Assessors (Li et al., 2024b; Zhang et al., 2024d), Critics (Ke et al., 2024; Xiong et al., 2024; Putta et al., 2024), Verifiers (Ling et al., 2024; Shinn et al., 2023; Wang et al., 2024e), Examiners (Bai et al., 2023a), Reward/Ranking Models (Yang et al., 2024b; Sun et al., 2023a; Luo et al., 2023a; Yuan et al., 2023), etc.
@@ 2.1.1. Generating scores
Figure 4. The template for Likert scale scoring from Gao et al. (2023b).
@@ 2.1.3. Conducting pairwise comparisons
Pairwise comparison is a well-established method that has significantly impacted a variety of fields (Qin et al., 2024a). As noted by (Liu et al., 2024b), LLM and human evaluations are more aligned in the context of pairwise comparisons compared to score-based assessments. Numerous studies have demonstrated that pairwise comparative assessments outperform other judging methods in terms of positional consistency (Zheng et al., 2023c; Liusie et al., 2024). Furthermore, pairwise comparisons can be extended to more complex relation-based assessment frameworks, such as list-wise comparisons, using advanced ranking algorithms (Qin et al., 2024a; Liu et al., 2024b), data filtering (Yuan et al., 2023). In pairwise comparative assessments, LLM-as-a-Judge is prompted to select the response that better answers the question at hand. To accommodate the possibility of a tie, several option modes are introduced. The Two-Option mode requires judges to choose the better response from two given options. The Three-Option mode introduces an additional choice, allowing judges to indicate a tie if neither response is preferable. Evaluations typically involve determining the outcomes of win, tie, or loss for responses (Wang et al., 2023d) through pairwise comparisons, with win rounds counted for each response. The Four-Option mode further expands the choices, allowing judges to classify responses as either a ”both good tie” or a ”both bad tie.”
@@
Figure 6. The template for pairwise comparison from Gao et al. (2023b)
@@ 2.2.2. Fine-tuned LLM
However, relying on external API for evaluation may introduce consideration about privacy leakage, and the opacity of API models also challenges the evaluation reproducibility. Therefore, subsequent studies recommend refining language models tailored for evaluations by emphasizing the use of pairwise comparisons or grading. For instance, PandaLM (Wang et al., 2023d) constructs data based on Alpaca instructions and GPT-3.5 annotation, and then fine-tunes LLaMA-7B (Touvron et al., 2023a) as an evaluator model. JudgeLM (Zhu et al., 2023a) constructs data from diversified instruction sets and GPT-4 annotations, and fine-tunes Vicuna (Touvron et al., 2023b) as a scalable evaluator model. Auto-J (Li et al., 2023b) constructs evaluation data upon multiple scenarios to train a generative evaluator model, which can provide both evaluation and critical opinion. Prometheus (Kim et al., 2023) defines thousands of evaluation criteria and constructs a feedback dataset based on GPT-4, and fine-tunes a fine-grained evaluator model.
The typical process for fine-tuning a judge model involves three main steps as shown in Figure 14. Step 1: Data Collection. The training data generally consists of three components: instructions, the objects to be evaluated, and evaluations. Instructions are typically sourced from instruction datasets, while evaluations can come from either GPT-4 or human annotations. Step 2-Prompt Design. The structure of the prompt template can vary based on the evaluation scheme, which already detailed in § 2.1. Step 3: Model Fine-Tuning. Using the designed prompts and collected data, the fine-tuning process for the evaluator model typically adheres to the instruction fine-tuning paradigm (Ouyang et al., 2022b). The model receives an instruction along with one or more responses to generate output that includes evaluation results and possibly explanations.
@@ 2.3.1. Extracting specific tokens
2.3.2. Constrained decoding
Constrained decoding is a technique that enforces structured output from Large Language Models (LLMs) by restricting token generation according to predefined schemas, typically in formats like JSON. This approach uses a finite state machine (FSM) to compute valid next tokens at each decoding step, effectively masking the model’s output probability distribution to ensure conformity with the desired schema. While this method guarantees syntactically valid outputs, it presents several challenges: it can distort the model’s learned distribution and potentially degrade output quality, requires significant engineering implementation effort, and introduces computational overhead during inference.
Recent work has proposed various solutions to address these challenges. (Beurer-Kellner et al., 2024) introduce DOMINO, a decoding algorithm that preserves natural tokenization while enforcing constraints. Their system minimizes overhead through precomputation and speculative decoding, sometimes achieving faster performance than unconstrained decoding. (Dong et al., 2024b) develop XGrammar, which accelerates grammar-constrained generation by separating tokens into those that can be pre-checked and those requiring runtime verification. By co-designing the grammar engine with LLM inference, they achieve up to 100x speedup over existing approaches.(Zheng et al., 2024b) present SGLang, combining a domain-specific language with an optimized runtime. Their system features efficient KV cache reuse and compressed finite state machines for faster decoding, demonstrating that thoughtful co-design of programming model and runtime can minimize constrained decoding overhead.
@@ 2.3.3. Normalizing the output logits
In addition, Self-evaluation (Hao et al., 2023a) is also common using this method for LLM-as-a-Judge. It can be helpful to let the LLM evaluate itself by asking, ”Is this reasoning step correct?” and then reward it based on the probability of the next word being ”Yes.”
@@ 2.3.4. Selecting sentences
In addition to selecting specific tokens and normalizing the output logits, the content extracted by LLM-as-a-Judge may also be a sentence or paragraph. As showed in Figure 2, agent for reasoning task (Hao et al., 2023a), builds a reasoning tree by iteratively considering the most promising reasoning steps (actions, sub-questions) by LLM-as-a-Judge.
@@ 2.4.1. LLM-as-a-Judge for Models
It is universally known that the best way to evaluate LLMs is human judgment, but collecting human annotations can be costly, time-consuming, and laborious (Ouyang et al., 2022b; Zheng et al., 2023d). Using strong LLMs (usually closed-source ones, e.g., GPT-4, Claude, ChatGPT) as an automated proxy for assessing LLMs has become a natural choice (Zhou et al., 2023a). With appropriate prompt design, the quality of evaluation and agreement to human judgment can be promising (Dubois et al., 2023; Zheng et al., 2023d; Zhang et al., 2023c; Wang et al., 2023b). However, the cost concern still exists when calling the APIs of these proprietary models, especially when there is a frequent need for model validation on large-scale data. Moreover, closed-source LLM-as-a-Judge leads to low reproducibility due to potential changes in models behind the API. Some recent works have started to make attempts for open-source alternatives. SelFee (Ye et al., 2023) collects generations, feedback, and revised generations from ChatGPT and fine-tunes LLaMA models to build a critique model. Shepherd (Wang et al., 2023c) trains a model that can output critiques for single-response with the data of feedback from online communities and human annotation. PandaLM (Wang et al., 2023d) trains a model to conduct pairwise comparison for LLM Instruction Tuning Optimization, and Zheng et al. (2023d) also fine-tune Vicuna (Touvron et al., 2023b) on a 20K pairwise comparison dataset to explore the potential of open-source models as a more cost-friendly proxy.
Figure 9. The illustrations of the scenario LLM-as-a-Judge for Models
@@ 2.4.2. LLM-as-a-Judge for Data
However, solely relying on LLM-as-a-Judge for data annotation poses challenges, particularly as the value of annotated data diminishes with the rapid improvement of model performance. To address this, approaches like Self-Taught Evaluator (Wang et al., 2024b) offer a promising alternative by eliminating the need for human annotations. This method leverages synthetic training data, starting with unlabeled instructions and generating contrasting outputs from models. These outputs are then used to train an LLM-as-a-Judge to produce reasoning traces and final judgments. With each iteration, the evaluator improves by learning from its refined predictions, creating a cycle of continuous self-enhancement. This iterative approach not only keeps annotations relevant but also ensures that evaluators evolve alongside advancing models.
@@ 2.4.3. LLM-as-a-Judge for Agents
There are two ways to apply LLM-as-a-Judge for an agent. One is to evaluate the entire process of the intelligent agent (Zhuge et al., 2024), and the other is to evaluate it at a specific stage in the agent framework process (Hao et al., 2023a; Shinn et al., 2023). Both approaches are briefly illustrated in Figure 10. Using LLM as the brain of agent, an agentic system (Zhuge et al., 2024) could evaluate like a human, it would reduce the need for human involvement and eliminate the trade-off between thoroughness and effort. In addition, the agent (Shinn et al., 2023) can interact with the environment through language and receive feedback on actions through LLM to make decisions for the next action.